Consider a population consisting of 1,2,3,4 and 5. From the population, the mean of the sampling distribution of the sample means if n=2, n=3. and n=4


1
Expert's answer
2021-03-23T05:10:36-0400

The mean of the sample means is

When n=2,

The samples are






The mean of sample mean "=\dfrac{\sum \bar{x}}{9}=\dfrac{28}{9}=3.11"


When n=3,

The samples are







The mean of sample mean "=\dfrac{\sum \bar{x}}{9}=\dfrac{26.1}{9}=2.9"


When n=4,

The samples are



The mean of sample mean "=\dfrac{\sum \bar{x}}{5}=\dfrac{17}{5}=3.4"